Which expression is equivalent to

(sin(θ)+cos(θ))^2?
1
Expert's answer
2020-11-17T16:16:54-0500

"(sin\varTheta +cos\varTheta)^2 =\\
 = sin^2\varTheta +2sin\varTheta*cos\varTheta +cos^2\varTheta=\\
= sin^2\varTheta+cos^2\varTheta +sin2\varTheta = \\
=1 + sin2\varTheta"

answer: "1 + sin2\varTheta"